OpenStudy (anonymous):

please help. Mark needs $26,700 to purchase a boat. How much money will he need to invest today in a savings account earning 4.6% interest, compounding quarterly, to have enough money to purchase the boat in 12 years? $3,083.12 $15,422.28 $22,220.50 $25,266.90

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

Using the compound interest formula, \[26,700=P\left(1+\frac{.046}{4}\right)^{4\times12}\]

OpenStudy (anonymous):

Then solve for \(P\)

OpenStudy (anonymous):

Ok so would I do what is in the parenthesis first?

OpenStudy (anonymous):

Yes, simplify the red part first, then divide both sides by what you get. \[26,700=P\color{red}{\left(1+\frac{.046}{4}\right)^{4\times12}}\]

OpenStudy (anonymous):

i cant get an answer i have been trying. @sithsandgiggles

OpenStudy (anonymous):

@ankit042

OpenStudy (anonymous):

@NaCl

OpenStudy (anonymous):

@AwesomeB

OpenStudy (ankit042):

@kenbaby Trying solving the equation you will get your answer

OpenStudy (anonymous):

@ankit042 i did try, after i solve whats in the parenthesis what is my next step

OpenStudy (ankit042):

P is the variable in the equation solve the parenthesis P=26700/(value in parenthesis)

OpenStudy (anonymous):

For some reason I am not getting the right answer.

OpenStudy (anonymous):

@ankit042

OpenStudy (ankit042):

can you tell the value of parenthesis?

OpenStudy (anonymous):

\[26,700=P\color{red}{\left(1+\frac{.046}{4}\right)^{4\times12}}\\ 26,700=P\color{red}{\left(1.0115\right)^{48}}\\ P=\frac{26,700}{\left(1.0115\right)^{48}}=\frac{26,700}{1.73126}=\cdots\]

OpenStudy (anonymous):

@sithsandgiggles im confused

OpenStudy (anonymous):

@ankit042 can you help